OpenStudy (anonymous):

1+sec^2x=tan^2x Is this an identity?

OpenStudy (anonymous):

no sec^2x=1+tan^2x

OpenStudy (anonymous):

yeahh from sin^2x + cos^2x = 1 and if divide everything by cos^2x we get tan^2x + 1 = sec^2x :)

OpenStudy (anonymous):

so therefore in what you wrong was wrong